# Open Heavy flavour mesons in hot asymmetric strange hadronic matter – a QCD sum rule approach

### v1

@@ Abstract
The in-medium masses of the pseudoscalar open charm ($D$, $\bar{D}$, $D_{s}$ and $\bar{D_{s}}$) and open bottom ($B$, $\bar{B}$, $B_{s}$ and $\bar{B_{s}}$) mesons in hot asymmetric strange hadronic matter are studied within a QCD sum rule approach. These are computed using the medium modifications of the light quark condensates ($\langle{\bar{q_{i}}}{q_{i}}\rangle$, with $(q_{i},i=1,2,3)\equiv(u,d,s)$), the scalar gluon condensate, $\left\langle\frac{\alpha_{s}}{\pi}G^{a}_{\mu\nu}{G^{a}}^{\mu\nu}\right\rangle$, the twist-2 tensorial gluon operator, $\left\langle\frac{\alpha_{s}}{\pi}\Big{(}G^{a}_{\mu\sigma}{{G^{a}}_{\nu}}^{\sigma}u^{\mu}u^{\nu}-\frac{1}{4}G^{a}_{\mu\sigma}{G^{a}}^{\mu\sigma}\Big{)}\right\rangle$ (where $u^{\mu}$ is the 4-velocity of the medium) and other operators in the operator product expansion upto mass dimension 5. Within a chiral SU(3) model, the quark condensates are obtained from the medium modifications of the non-strange and strange scalar-isoscalar fields ($\sigma$ and $\zeta$), and the scalar-isovector field, $\delta$, whereas, the gluon (scalar and twist-2) condensates are computed from the in-medium value of the dilaton field, $\chi$, which is incorporated within the chiral model to mimic the broken scale invariance of QCD. The mixed quark-gluon condensate $\langle\bar{q_{i}}g_{s}\sigma.Gq_{i}\rangle$ is calculated from the in-medium quark condensate $\langle\bar{q_{i}}q_{i}\rangle$. The splittings of the masses of the $D-\bar{D}$ ($B-\bar{B})$, as well as $D_{s}-\bar{D_{s}}(B_{s}-\bar{B_{s}})$ in the hadronic medium are due to the odd part of the spectral function. For the strange-charm (strange-bottom) mesons ($(q_{i},i=3)\equiv s$), the value of $\langle{q_{i}}^{\dagger}{q_{i}}\rangle$ dominates over the ocntributions from the other operators, $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$ of the odd part of the sectral function. This is observed as opposite behaviour of the particle-antiparticle mass splittings with the mass of the particle to be larger (smaller) than the antiparticle mass in nuclear (hyperonic) matter. The effects of density, isospin asymmetry, strangeness and temperature on the masses of the open charm and open bottom mesons are observed to be appreciable.
@@ I Introduction
The study of the properties of hadrons under extreme conditions of temperature and/or density is an important subject of research in strong interaction physics. The topic is of direct relevance for the experimental observables of the relativistic heavy-ion collision experiments, as well as, is relevant for study of the bulk properties of compact objects, e.g., the neutron stars. The light vector mesons have been extensively studied in the literature, as the medium modifications of these mesons can be observed in the dilepton spectra arising from heavy ion collision experiments Rapp_Wambach . In the recent years, a lot of studies on the in-medium properties of heavy flavour hadrons have been initiated due to its relevance and possible observability in the relativistic heavy ion collision experiments Prog_Nucl_Part_Phys_2017_Hosaka_et_al . The spectral modification of the charm hadrons will be investigated via $\bar{p}$-A collisions at FAIR at the future facility at GSI, Darmstadt Yamagata-Sekihara_et_al ; Rapp and charm production using pion beams on proton in J-PARC Kim1 ; Kim2 ; Garzon . The charm (bottom) meson-nucleon interaction can also lead to possibility of formation of charm (bottom) mesons bound in atomic-nuclei.
@@
The properties of heavy flavour mesons in the (hot) hadronic matter have been studied extensively in the literature. These have been investigated using the potential models potential_models , the coupled channel approach coupled_channel_approach , the quark meson coupling (QMC) model QMC_models ; krein_jpsi ; N_Quarkonia_krein heavy quark symmetry and interaction of these mesons with nucleons via pion exchange Yasui_Sudoh_pion , heavy meson effective theory Yasui_Sudoh_heavy_meson_Eff_th , studying the heavy flavour meson as an impurity in nuclear matter Yasui_Sudoh_heavy_particle_impurity , and the QCD sum rule approach open_heavy_flavour_qsr_1 ; kimlee ; klingl ; amarvjpsi_qsr ; moritalee_1 ; moritalee_2 ; moritalee_3 ; moritalee_4 ; leemorita2 ; open_heavy_flavour_qsr_2 ; open_heavy_flavour_qsr_3 ; open_heavy_flavour_qsr_4 ; Wang_heavy_mesons_1 ; Wang_heavy_mesons_2 ; arvind_heavy_mesons_QSR_1 ; arvind_heavy_mesons_QSR_2 ; arvind_heavy_mesons_QSR_3 . Using leading order QCD formula pes1 ; pes2 ; voloshin , the medium modifications of the masses of the charmonium states due to the medium change of the scalar gluon condensate were studied in a linear density approximation in Ref.leeko , which showed much larger mass shifts for the excited states ($\psi(3686)$ and $\psi(3770)$), as compared to the mass shift of $J/\psi$. Within a chiral effective model sche1 ; heide1 ; paper3 ; kristof1 ; hartree , the in-medium masses of the charmonium (bottomonium) states amarvdmesonTprc ; amarvepja ; AM_DP_upsilon as well as of the open heavy flavor mesons amdmeson ; amarindamprc ; amarvdmesonTprc ; amarvepja ; DP_AM_Ds ; DP_AM_bbar ; DP_AM_Bs have been studied. Within the chiral effective model, the mass modifications of the quarkonium states arise due to the medium modification of a scalar dilaton field, which is introduced in the model to mimic the broken scale invariance of QCD. The open heavy flavour mesons have been studied due to their interactions with the baryons and the scalar mesons in the hadronic medium. The model has been used to study the partial decay widths of the heavy quarkonium states to the open heavy flavour mesons, in the (magnetized) hadronic medium amarvepja ; charmdecay_mag using a light quark pair creation model friman , namely the ${}^{3}P_{0}$ model 3p0_1 ; 3p0_2 ; 3p0_3 ; 3p0_4 as well as using a field theoretical model for composite hadrons amspmwg ; amspm_upsilon ; charmdw_mag ; upslndw_mag . Within the chiral effective model, in the presence of an external magnetic field, the effects of magnetized Dirac sea have been investigated on the properties of heavy flavour mesons in the hadronic medium HQ_DW_DS_AM_SPM_2023 ; Charmonium_DS ; Open_Bottom_DS The magnetized Dirac sea, accounting for the anomalous magnetic moment of nucleons is observed to lead to the phenomenon of inverse magnetic catalysis at finite densities for symmetric as well as asymmetric nuclear matter in presence of strong magnetic fields. Recently, the effects of the Dirac sea as well as mixing of the pseudoscalar and vector mesons on the masses, decay widths and the production cross-sections of the charmonium states in hot magnetized strange hadronic matter have been investigated in Ref. AMAKSPM24 .
@@
Within the chiral SU(3) model, using the medium changes of the light quark condensates ($\langle\bar{u}u\rangle$, $\langle\bar{d}d\rangle$ and $\langle\bar{s}s\rangle$), obtained from the non-strange isoscalar ($\sigma$) isovector ($\delta$) and strange ($\zeta$) scalar fields and gluon condensates calculated from the scalar dilaton field, $\chi$, the light vector mesons ($\omega$, $\rho$ and $\phi$), in (magnetized) hadronic matter have been studied within the framework of QCD sum rule approach am_vecmeson_qsr ; vecqsr_mag . The effects of the magnetized Dirac sea on the spectral properties of light vector and axial-vector mesons in hot magnetized nuclear matter Spectr_rho_A1_DS and of the ground state heavy quarkonia QCDSR_Heavy_Quarkonia have also been studied with ths QCD sum rule approach. In the present work, we investigate the masses of the open charm ($D$, $\bar{D}$, $D_{s}$) and open bottom ($B$, $\bar{B}$, $B_{s}$) mesons in isospin asymmetric strange hadronic matter at finite temperature. In section II, we briefly describe the QCD sum rule approach used to compute the masses of the open charm and open bottom mesons, retaining the operators upto dimension 5 in the operator product expansion of the spectral function. Unlike the case of vacuum, in the hadronic medium, there is splitting between the masses of the particles and their antiparticles arising due to the odd part of the spectral function. In section III, we describe the chiral SU(3) model used to compute the in-medium values of the quark (light and strange), the gluon (scalar and twist-2) and the mixed quark-gluon operators, which are used for the calculation of the in-medium masses of the pseudoscalar open charm and open bottom mesons. Section IV discusses the results and in section V, we summarize the findings of the present work.
@@ II QCD sum rule approach
The QCD sum rule approach relates the hadronic properties, encoded in spectral functions, to the QCD condensates Cohen_Prog_Part_Nucl_Phys . The basic ingredient is the two-point correlation function given as
$\Pi(q)=i\int d^{3}xe^{-iq.x}\langle\Omega|T(j(x)j(0)^{\dagger})|\Omega\rangle,$ (1)
where, $|\Omega\rangle$ is the physical ground state. Unlike the case of vacuum, the two-point correction function in the hadronic medium has contributions which are odd in $q_{0}$. This can be written in terms of even and odd functions of $q_{0}$ as
@@
with $\Pi^{e}(q_{0},\vec{q})=\frac{1}{2}(\Pi(q_{0},\vec{q})+\Pi(-q_{0},\vec{q}))$ and $\Pi^{o}(q_{0},\vec{q})=\frac{1}{2q_{0}}(\Pi(q_{0},\vec{q})-\Pi(-q_{0},\vec{q}))$. It might be noted here that $\Pi^{e(o)}(q_{0},\vec{q})$ depends on $q_{0}^{2}$. The Borel transformed sum rules for the even and odd parts are given as open_heavy_flavour_qsr_2
$\Pi^{e}_{OPE}(M^{2})=\Bigg{[}\int_{-\infty}^{s_{0}^{-}}+\int_{s_{0}^{-}}^{s_{0}^{+}}+\int_{s_{0}^{+}}^{\infty}\Bigg{]}\frac{s\;{\rm d}s\;{\rm Im}\;\Pi(s)}{\pi}e^{-{s}/{M^{2}}}$ (3)
@@
$\Pi^{o}_{OPE}(M^{2})=\Bigg{[}\int_{-\infty}^{s_{0}^{-}}+\int_{s_{0}^{-}}^{s_{0}^{+}}+\int_{s_{0}^{+}}^{\infty}\Bigg{]}\frac{{\rm d}s\;{\rm Im}\;\Pi(s)}{\pi}e^{-{s}/{M^{2}}}.$ (4)
In the above equations, the subscript ‘$OPE$’ denotes the operator product exapnsion $\langle\Omega|T(j(x)j(0)^{\dagger})|\Omega\rangle=\sum_{n}C_{n}(x-y)\langle\Omega|O_{n}|\Omega\rangle$, with $\langle\Omega|O_{n}|\Omega\rangle$ as the QCD condensates for the operators $O_{n}$ and $C_{n}(x-y)$ are the Wilson coefficients, $s_{0}^{\pm}$ are the continuum thresholds, for the particle (antiparticle) and $M$ is the Borel mass. The open charm and open bottom mesons have been studied in symmetric nuclear matter within the QCD sum rue approach keeping operators upto dimension 5 in the OPE side of the spectral functions in Ref. open_heavy_flavour_qsr_2 . In the present work, we study the in-medium masses of these pseudoscalar mesons in hot isospin asymmetric strange hadronic matter. These spectral functions are obtained using the current operators for the particles ($p_{i}$), $j_{i}=j^{p}_{i}\equiv i{\bar{q_{i}}}\gamma_{5}Q$, with $(q_{i},i=1,2,3)\equiv(u,d,s)$, where, the heavy quark $Q=c$ corresponds to $D^{0}$, $D^{+}$ and $D_{s}^{+}$ mesons, and, $Q=b$ corresponds to $B^{-}$, $\bar{B^{0}}$ and $\bar{B_{s}}^{0}$ mesons respectively. The current operator is $j_{i}^{\dagger}\equiv i{\bar{Q}}\gamma_{5}{q_{i}}$, for the antiparticles ($\bar{D^{0}}$, $D^{-}$, $D_{s}^{-}$ for charm sector and $B^{+}$, $B^{0}$ and ${B_{s}}^{0}$ for the bottom sector). The Borel transformed even and odd parts of the OPE upto dimension 5 are obtained as open_heavy_flavour_qsr_2
$\Pi^{e}_{OPE}(M^{2})=\frac{1}{\pi}\int_{m_{Q}^{2}}^{\infty}{\rm Im}\Pi^{per}_{P}(s,\vec{q}=0)+e^{-{m_{Q}^{2}}/{M^{2}}}P^{e}(M^{2})$ (5)
@@
$+$ $4\Big{(}\frac{m_{Q}^{3}}{2M^{4}}-\frac{m_{Q}}{M^{2}}\Big{)}\Big{[}\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle-\frac{1}{8}\langle\bar{q_{i}}g_{s}\sigma\cdot Gq_{i}\rangle\Big{]}$
@@
In equation (6), the first and second terms arise due to the quark-antiquark condensate ($\langle{\bar{q_{i}}}q_{i}\rangle$), retaining the term proportional to the light quark mass ($m_{q_{i}}$) open_heavy_flavour_qsr_1 as given by the second term. The first term, due to the large mass of the heavy quark ($m_{Q}$) is the dominant contribution to the mass of the open charm (bottom) meson, arising from the dimension 3 quark-antiquark condensate in the operator product exapansion. The strong coupling constant, $\alpha_{s}$ is given as $\alpha_{s}=4\pi/\Big{[}\big{(}(11-2N_{f}/3)\rm{ln}(\mu^{2}/\Lambda_{QCD}^{2})\big{)}\Big{]}$, with $\Lambda_{QCD}=500$ MeV, and, $N_{f}$=4 (5) for the study of the open charm (bottom) mesons. $\mu$, the renormalization scale, is taken to be mass of the largest quark in the system, i.e, the mass of the charm (bottom) quark, for investigating the mass of the open charm (bottom) mesons open_heavy_flavour_qsr_2 , and $\gamma_{E}$ is the Euler constant. The other terms are the higher mass dimension operators, which include the dimension-4 scalar and twist-2 gluon condensate, the mixed quark-gluon operator (the abbreviation $\sigma\cdot G$ is used for $\sigma_{\mu\nu}G^{\mu\nu}$), and terms proportional to the operators $\langle{q_{i}}^{\dagger}iD_{0}q_{i}\rangle$ and $\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle$. In the odd part of the Borel transformed spectral function, the dominant contribution in eqn (8) is from the number density of the quark ($\langle{q_{i}}^{\dagger}{q_{i}}\rangle$), along with the contributions from the dimension 5 operators, $\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma\cdot Gq_{i}\rangle$. In the term proportional to the twist-2 gluon condensate operator, $u^{\mu}$ is the 4-velocity of the hadronic medium. For the spectral function, the ansatz
${\rm Im}\Pi(s)=\pi F_{+}\delta(s-m_{+})-\pi F_{-}\delta(s+m_{-}),$ (9)
is used, where, $m_{\pm}$ correspond to the mass of the particle (antiparticle). We choose the continuum thresholds as $s_{0}^{\pm}=\pm s_{0}$, similar to the case of vacuum. The masses for the particles and antiparticle, $m_{\pm}$, corresponding to the open heavy flavour mesons, are solved by using the spectral function given by equation (9) in the equations (3) and (4) open_heavy_flavour_qsr_2 . The expressions for the masses of the particle and antiparticle are given as where,

### v2

# Open Heavy flavor mesons in hot asymmetric strange hadronic matter – a QCD sum rule approach

@@ Abstract
The in-medium masses of the pseudoscalar open charm ($D$, $\bar{D}$, $D_{s}$ and $\bar{D_{s}}$) and open bottom ($B$, $\bar{B}$, $B_{s}$ and $\bar{B_{s}}$) mesons in hot asymmetric strange hadronic matter are studied within a QCD sum rule approach. These are computed using the medium modifications of the light quark condensates ($\langle{\bar{q_{i}}}{q_{i}}\rangle$, with $(q_{i},i=1,2,3)\equiv(u,d,s)$), the scalar gluon condensate, $\left\langle\frac{\alpha_{s}}{\pi}G^{a}_{\mu\nu}{G^{a}}^{\mu\nu}\right\rangle$, the twist-2 tensorial gluon operator, $\left\langle\frac{\alpha_{s}}{\pi}\Big{(}G^{a}_{\mu\sigma}{{G^{a}}_{\nu}}^{\sigma}u^{\mu}u^{\nu}-\frac{1}{4}G^{a}_{\mu\sigma}{G^{a}}^{\mu\sigma}\Big{)}\right\rangle$ (where $u^{\mu}$ is the 4-velocity of the medium) and other operators in the operator product expansion upto mass dimension 5. Within a chiral SU(3) model, the quark condensates are obtained from the medium modifications of the non-strange and strange scalar-isoscalar fields ($\sigma$ and $\zeta$), and the scalar-isovector field, $\delta$, whereas, the gluon (scalar and twist-2) condensates are computed from the in-medium value of the dilaton field, $\chi$, which is incorporated within the chiral model to mimic the broken scale invariance of QCD. The mixed quark-gluon condensate $\langle\bar{q_{i}}g_{s}\sigma.Gq_{i}\rangle$ is calculated from the in-medium quark condensate $\langle\bar{q_{i}}q_{i}\rangle$. The splittings of the masses of the $D-\bar{D}$ ($B-\bar{B})$, as well as $D_{s}-\bar{D_{s}}(B_{s}-\bar{B_{s}})$ in the hadronic medium are due to the odd part of the spectral function. For the strange-charm (strange-bottom) mesons ($(q_{i},i=3)\equiv s$), the value of $\langle{q_{i}}^{\dagger}{q_{i}}\rangle$ dominates over the contributions from the other operators, $\langle{q_{i}}^{\dagger}iD_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma.G{q_{i}}\rangle$, of the odd part of the sectral function. This is observed as opposite behaviour of the particle-antiparticle mass splittings with the mass of the particle to be larger (smaller) than the antiparticle mass in nuclear (hyperonic) matter. The density and isospin asymmetry effects are observed to be the dominant medium effects which might have observable consequences on the production of charmonia and open charm mesons, and, on the production ratios, $D^{+}/D^{-}$ and $D^{0}/\bar{D^{0}}$, in asymmetric heavy ion collisions in the CBM experiment at FAIR at the future facility at GSI.
@@ I Introduction
The study of the properties of hadrons under extreme conditions of temperature and/or density is an important subject of research in strong interaction physics. The topic is of direct relevance for the experimental observables of the relativistic heavy-ion collision experiments, as well as, is relevant for study of the bulk properties of compact objects, e.g., the neutron stars. The light vector mesons have been extensively studied in the literature, as the medium modifications of these mesons can be observed in the dilepton spectra arising from heavy ion collision experiments Rapp_Wambach . In the recent years, a lot of studies on the in-medium properties of heavy flavor hadrons have been initiated due to its relevance and possible observability in the relativistic heavy ion collision experiments Prog_Nucl_Part_Phys_2017_Hosaka_et_al . The spectral modification of the charm hadrons will be investigated via $\bar{p}$-A collisions at FAIR at the future facility at GSI, Darmstadt Yamagata-Sekihara_et_al ; Rapp and charm production using pion beams on proton in Japan Proton Accelerator Research Complex (J-PARC) Kim1 ; Kim2 ; Garzon . The charm (bottom) meson-nucleon interaction can also lead to possibility of formation of charm (bottom) mesons bound in atomic nuclei.
@@
The properties of heavy flavor mesons in the (hot) hadronic matter have been studied extensively in the literature. These have been investigated using the potential models potential_models , the coupled channel approach coupled_channel_approach , the quark meson coupling (QMC) model QMC_models ; krein_jpsi ; N_Quarkonia_krein heavy quark symmetry and interaction of these mesons with nucleons via pion exchange Yasui_Sudoh_pion , heavy meson effective theory Yasui_Sudoh_heavy_meson_Eff_th , studying the heavy flavor meson as an impurity in nuclear matter Yasui_Sudoh_heavy_particle_impurity , and the QCD sum rule approach open_heavy_flavour_qsr_1 ; kimlee ; klingl ; amarvjpsi_qsr ; moritalee_1 ; moritalee_2 ; moritalee_3 ; moritalee_4 ; leemorita2 ; open_heavy_flavour_qsr_2 ; Hilger_thesis ; open_heavy_flavour_qsr_3 ; open_heavy_flavour_qsr_4 ; Wang_heavy_mesons_1 ; Wang_heavy_mesons_2 ; arvind_heavy_mesons_QSR_1 ; arvind_heavy_mesons_QSR_2 ; arvind_heavy_mesons_QSR_3 . Using leading order QCD formula pes1 ; pes2 ; voloshin , the medium modifications of the masses of the charmonium states due to the medium change of the scalar gluon condensate were studied in a linear density approximation in Ref.leeko , which showed much larger mass shifts for the excited states ($\psi(3686)$ and $\psi(3770)$), as compared to the mass shift of $J/\psi$. Within a chiral effective model sche1 ; heide1 ; paper3 ; kristof1 ; hartree , the in-medium masses of the charmonium (bottomonium) states amarvdmesonTprc ; amarvepja ; AM_DP_upsilon as well as of the open heavy flavor mesons amdmeson ; amarindamprc ; amarvdmesonTprc ; amarvepja ; DP_AM_Ds ; DP_AM_bbar ; DP_AM_Bs have been studied. Within the chiral effective model, the mass modifications of the quarkonium states arise due to the medium modification of a scalar dilaton field, which is introduced in the model to mimic the broken scale invariance of QCD. The open heavy flavor mesons have been studied due to their interactions with the baryons and the scalar mesons in the hadronic medium. The model has been used to study the partial decay widths of the heavy quarkonium states to the open heavy flavor mesons, in the (magnetized) hadronic medium amarvepja ; charmdecay_mag using a light quark pair creation model friman , namely the ${}^{3}P_{0}$ model 3p0_1 ; 3p0_2 ; 3p0_3 ; 3p0_4 as well as using a field theoretical model for composite hadrons amspmwg ; amspm_upsilon ; charmdw_mag ; upslndw_mag . Within the chiral effective model, in the presence of an external magnetic field, the effects of magnetized Dirac sea have been investigated on the properties of heavy flavor mesons in the hadronic medium HQ_DW_DS_AM_SPM_2023 ; Charmonium_DS ; Open_Bottom_DS . The magnetized Dirac sea, accounting for the anomalous magnetic moment of nucleons is observed to lead to the phenomenon of inverse magnetic catalysis at finite densities for symmetric as well as asymmetric nuclear matter in presence of strong magnetic fields. Recently, the effects of the Dirac sea as well as mixing of the pseudoscalar and vector mesons on the masses, decay widths and the production cross-sections of the charmonium states in hot magnetized strange hadronic matter have been investigated in Ref. AMAKSPM24 .
@@
Within the chiral SU(3) model, using the medium changes of the light quark condensates ($\langle\bar{u}u\rangle$, $\langle\bar{d}d\rangle$ and $\langle\bar{s}s\rangle$), obtained from the non-strange isoscalar ($\sigma$) isovector ($\delta$) and strange ($\zeta$) scalar fields and gluon condensates calculated from the scalar dilaton field, $\chi$, the light vector mesons ($\omega$, $\rho$ and $\phi$), in (magnetized) hadronic matter have been studied within the framework of QCD sum rule approach am_vecmeson_qsr ; vecqsr_mag . The effects of the magnetized Dirac sea on the spectral properties of light vector and axial-vector mesons in hot magnetized nuclear matter Spectr_rho_A1_DS and of the ground state heavy quarkonia QCDSR_Heavy_Quarkonia have also been studied with the QCD sum rule approach. In the present work, we investigate the masses of the open charm ($D$, $\bar{D}$, $D_{s}$) and open bottom ($B$, $\bar{B}$, $B_{s}$) mesons in isospin asymmetric strange hadronic matter at finite temperature. In section II, we briefly describe the QCD sum rule approach used to compute the masses of the open charm and open bottom mesons, retaining the operators upto dimension 5 in the operator product expansion of the spectral function. Unlike the case of vacuum, in the hadronic medium, there is splitting between the masses of the particles and their antiparticles arising due to the odd part of the spectral function. In section III, we describe the chiral SU(3) model used to compute the in-medium values of the quark (light and strange), the gluon (scalar and twist-2) and the mixed quark-gluon operators, which are used for the calculation of the in-medium masses of the pseudoscalar open charm and open bottom mesons. Section IV discusses the results and in section V, we summarize the findings of the present work.
@@ II QCD sum rule approach
The QCD sum rule approach relates the hadronic properties, encoded in spectral functions, to the QCD condensates Cohen_Prog_Part_Nucl_Phys ; QCD_Cond_Gubler_Satow_Prog_Part_Nucl_Phys . The basic ingredient is the two-point correlation function given as
$\Pi(q)=i\int d^{4}xe^{-iq.x}\langle\Omega|T(j(x)j(0)^{\dagger})|\Omega\rangle,$ (1)
where, $|\Omega\rangle$ is the physical ground state. Unlike the case of vacuum, the two-point correlation function in the hadronic medium has contributions which are odd in $q_{0}$, and, can be written in terms of even and odd functions of $q_{0}$ as
@@
with $\Pi^{e}(q_{0},\vec{q})=\frac{1}{2}(\Pi(q_{0},\vec{q})+\Pi(-q_{0},\vec{q}))$ and $\Pi^{o}(q_{0},\vec{q})=\frac{1}{2q_{0}}(\Pi(q_{0},\vec{q})-\Pi(-q_{0},\vec{q}))$. It might be noted here that $\Pi^{e(o)}(q_{0},\vec{q})$ depends on $q_{0}^{2}$. The Borel transformed sum rules for the even and odd parts, for $\vec{q}=0$, are given as open_heavy_flavour_qsr_2
${\cal B}[\Pi^{e}_{OPE}(\omega^{2},\vec{q}=0)](M^{2})=\frac{1}{\pi}\Big{[}\int_{s_{0}^{-}}^{s_{0}^{+}}{\rm d}s\;{\rm{Im}}\Pi^{(ph)}(s)+\Big{(}\int_{-\infty}^{s_{0}^{-}}+\int_{s_{0}^{+}}^{\infty}\Big{)}{\rm d}s\;{\rm{Im}}\Pi(s)\Big{]}s\;e^{-\frac{s^{2}}{M^{2}}},$ (3)
@@
${\cal B}[\Pi^{o}_{OPE}(\omega^{2},\vec{q}=0)](M^{2})=\frac{1}{\pi}\Big{[}\int_{s_{0}^{-}}^{s_{0}^{+}}{\rm d}s\;{\rm{Im}}\Pi^{(ph)}(s)+\Big{(}\int_{-\infty}^{s_{0}^{-}}+\int_{s_{0}^{+}}^{\infty}\Big{)}{\rm d}s\;{\rm{Im}}\Pi(s)\Big{]}e^{-\frac{s^{2}}{M^{2}}},$ (4)
In the above equations, $\omega^{2}=-q_{0}^{2}$, the subscript ‘$OPE$’ denotes the operator product expansion $\langle\Omega|T(j(x)j(y)^{\dagger})|\Omega\rangle=\sum_{n}C_{n}(x-y)\langle\Omega|O_{n}|\Omega\rangle$, where $\langle\Omega|O_{n}|\Omega\rangle$ are the QCD condensates for the operators $O_{n}$, $C_{n}(x-y)$ are the Wilson coefficients, $s_{0}^{+}(s_{0}^{-})$ is the continuum threshold for the particle (antiparticle) and $M$ is the Borel mass. The open charm and open bottom mesons have been studied in symmetric nuclear matter within the QCD sum rue approach keeping operators upto dimension 5 in the OPE side of the spectral function in Ref. open_heavy_flavour_qsr_2 . In the present work, we study the in-medium masses of these pseudoscalar mesons in hot isospin asymmetric strange hadronic matter, again, retaining QCD operators upto dimension 5 in the operator product expansion. The spectral functions of the open heavy flavor mesons are obtained using the current operators for the particles ($p_{i}$), $j_{i}=j^{p}_{i}\equiv i{\bar{q_{i}}}\gamma_{5}Q$, with $(q_{i},i=1,2,3)\equiv(u,d,s)$, where, the heavy quark $Q=c$ corresponds to $D^{0}$, $D^{+}$ and $D_{s}^{+}$ mesons, and, $Q=b$ corresponds to $B^{-}$, $\bar{B^{0}}$ and $\bar{B_{s}}^{0}$ mesons respectively. The current operator is $j_{i}^{\dagger}\equiv i{\bar{Q}}\gamma_{5}{q_{i}}$, for the antiparticles ($\bar{D^{0}}$, $D^{-}$, $D_{s}^{-}$ for charm sector and $B^{+}$, $B^{0}$ and ${B_{s}}^{0}$ for the bottom sector). For the phenomenological hadronic spectral function, we take the ansatz open_heavy_flavour_qsr_2
${\rm Im}\Pi^{(ph)}(s)=\pi F_{+}\delta(s-m_{+})-\pi F_{-}\delta(s+m_{-}),$ (5)
where $m_{\pm}$ corresponds to the mass of the particle (antiparticle). In equations (3) and (4), we change the integration variable from $s$ to $\tilde{s}\;(=s^{2})$ in the integrals corresponding to the continuum parts of the spectral functions, and, choose the continuum thresholds for the particle and antiparticle, to be $s_{0}^{\pm}=\pm\sqrt{{\tilde{s}}_{0}}$, similar to the case of vacuum. We also assume, according to semi-local duality hypothesis, that ${\rm Im}\Pi({\tilde{s}})$ corresponding to the particle and antiparticle continuum parts of the spectral function on the right hand side of the equations (3) and (4) to be same as the perturbative contribution of the spectral function, ${\rm Im}\Pi_{OPE}^{per}(s)$, as calculated in operator product expansion Hilger_thesis . It might be noted here that for the odd part of the spectral function, the perturbative contributions from the particle and antiparticle cancel with each other for the choice of the continuum thresholds $s_{0}^{\pm}=\pm\sqrt{{\tilde{s}}_{0}}$, as in the present work. The Borel transformed sum rules for the even and odd parts of the spectral function are then obtained as open_heavy_flavour_qsr_2
${\cal B}[\Pi^{e}_{OPE}(\omega^{2},\vec{q}=0)](M^{2})=\frac{1}{\pi}\int_{m_{Q}^{2}}^{\infty}{\rm d}{\tilde{s}}{\rm Im}\;\Pi^{per}_{OPE}({\tilde{s}},\vec{q}=0){e^{-{\tilde{s}}/M^{2}}}+e^{-{m_{Q}^{2}}/{M^{2}}}P^{e}(M^{2})$ (6)
$=$ $m_{+}F_{+}e^{-{m_{+}^{2}}/{M^{2}}}+m_{-}F_{-}e^{-{m_{-}^{2}}/{M^{2}}}+\frac{1}{\pi}\int_{{\tilde{s}}_{0}}^{\infty}{\rm d}{\tilde{s}}{\rm Im}\Pi^{per}_{OPE}({\tilde{s}})e^{-{\tilde{s}}/{M^{2}}}$
@@
$+$ $4\Big{(}\frac{m_{Q}^{3}}{2M^{4}}-\frac{m_{Q}}{M^{2}}\Big{)}\Big{[}\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle-\frac{1}{8}\langle\bar{q_{i}}g_{s}\sigma\cdot Gq_{i}\rangle\Big{]},$
@@
The QCD condensates corresponding to the even and odd parts of the spectral function occur in the polynomials $P^{e}(M^{2})$ and $P^{o}(M^{2})$, given by equations (7) and (9) respectively. For the polynomial $P^{e}(M^{2})$, the first and second terms of the arise due to the quark-antiquark condensate ($\langle{\bar{q_{i}}}q_{i}\rangle$), retaining the term proportional to the light quark mass ($m_{q_{i}}$) open_heavy_flavour_qsr_1 as given by the second term. The first term, due to the large mass of the heavy quark ($m_{Q}$) is the dominant contribution to the mass of the open charm (bottom) meson, arising from the dimension 3 quark-antiquark condensate in the operator product expansion. The strong coupling constant, $\alpha_{s}$ is given as $\alpha_{s}=4\pi/\Big{[}\big{(}(11-2N_{f}/3)\rm{ln}(\mu^{2}/\Lambda_{QCD}^{2})\big{)}\Big{]}$, with $\Lambda_{QCD}=500$ MeV, and, $N_{f}$=4 (5) for the study of the open charm (bottom) mesons. $\mu$, the renormalization scale, is taken to be mass of the largest quark in the system, i.e, the mass of the charm (bottom) quark, for investigating the mass of the open charm (bottom) mesons open_heavy_flavour_qsr_2 , $\gamma_{E}$ is the Euler constant. The other terms are the higher mass dimension operators, which include the dimension-4 scalar and twist-2 gluon condensates, the mixed quark-gluon operator (the abbreviation $\sigma\cdot G$ is used for $\sigma_{\mu\nu}G^{\mu\nu}$), and terms proportional to the operators $\langle{q_{i}}^{\dagger}iD_{0}q_{i}\rangle$ and $\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle$. In the term proportional to the twist-2 gluon condensate operator, $u^{\mu}$ is the 4-velocity of the hadronic medium. We study the spectral functions of the open heavy flavor mesons in the rest frame of the hadronic medium, for which $u^{\mu}=(1,\vec{u}=0$). In the odd part of the Borel transformed spectral function, the dominant contribution in equation (9) is from the number density of the quark ($\langle{q_{i}}^{\dagger}{q_{i}}\rangle$), along with the contributions from the dimension 5 operators, $\langle\bar{q_{i}}D_{0}^{2}{q_{i}}\rangle$ and $\langle{q_{i}}^{\dagger}g_{s}\sigma\cdot Gq_{i}\rangle$. The masses for the particles and antiparticle, $m_{\pm}$, corresponding to the open heavy flavor mesons, are solved by using the spectral function given by equation (5) in the equations (3) and (4) open_heavy_flavour_qsr_2 . The expressions for the masses of the particle and antiparticle are given as Hilger_thesis